OpenStudy (anonymous):

√[(1-sinA)/(1+sinA)]=secA-tanA

OpenStudy (callisto):

\[\sqrt{\frac{(1-sinA)}{(1+sinA)}} \]\[= \sqrt{\frac{(1-sinA)(1-sinA)}{(1+sinA)(1-sinA)}} \]\[=\sqrt{\frac{(1-sinA)^2}{(1-sin^2A)}} \]\[=\sqrt{\frac{(1-sinA)^2}{(cos^2A)}} \]\[=\frac{(1-sinA)}{(cosA)} \]\[=\frac{1}{cosA}-\frac{sinA}{cosA} \]\[=secA-tanA\]

OpenStudy (anonymous):

thanxx

OpenStudy (callisto):

Welcome.